# KodCode : A Diverse, Challenging, and Verifiable Synthetic Dataset for Coding

### v1

@@ 1 Introduction
Ideally, training high-performing coding LLMs requires high-quality data with verified solutions and test cases for post-training stages, including supervised fine-tuning (SFT) and reinforcement learning (RL) DeepSeek-AI et al. (2025); Team (2025); Hui et al. (2024); Wei et al. (2024a). While human-curated coding datasets like TACO Li et al. (2023), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022) offer high quality questions, canonical solutions, and tests, their limited scale constrains model training. Synthetic datasets have emerged as an alternative Wang et al. (2023); Long et al. (2024), but often lack diversity Xu et al. (2024), sufficient complexity Luo et al. (2023), and reliable response verification Lei et al. (2024).
@@
Furthermore, to validate KodCode’s effectiveness for code LLM post-training, we evaluate models fine-tuned on KodCode across standard benchmarks: HumanEval(+), MBPP(+), BigCodeBench, and LiveCodeBench. Our experimental results show that KodCode-fine-tuned models achieve state-of-the-art performance, surpassing other open-source models such as Qwen2.5-Coder-32B-Instruct and DeepSeek-R1-Distill-Llama-70B in most of the benchmarks.
@@ 2.2 Step 2: Solution & Test Generation
To generate verifiably correct coding solutions and unit tests $(sol,test)$ for questions from Step 1, we employ a self-verification procedure as detailed below. To ensure quality of solution and tests, we first employ GPT-4o-0513 (which achieves state-of-the-art performance among non-reasoning models on the BigCodeBench Leaderboard (Zhuo et al., 2024)) to generate both solution and test, then execute these unit tests to validate the correctness of the solution. Only question-solution-test triplets that pass self-verification are retained. The complete prompt template is provided in Appendix D.6.
Since our goal is to generate verifiable training data that is challenging and diverse in coverage; however, even state-of-the-art models cannot guarantee bug-free code and ensure solutions pass their unit tests. Simply discarding questions that fail self-verification risks eliminating many challenging ones, potentially biasing our dataset towards a distribution of simple problems.
Our solution approach to address this challenge is the allocation of additional self-verification attempts for hard questions. For each question from Step 1, we allow up to a maximum of $n$ attempts (where $n=10$ in our experiments) to generate a solution that passes its unit tests. Importantly, each attempt fixes the question and regenerates both the solution and its corresponding unit tests from scratch. This approach preserves challenging questions while naturally assigning difficulty labels based on the success rate across attempts. Questions that fail to generate correct solutions after $n$ attempts are discarded as they likely contain inherent flaws. Upon completing Step 2, we obtain a collection of 279K verified triplets.
@@ 2.3 Step 3: Post-training Data Synthesis
Motivated by recent advances in reasoning models (Team, 2025; Labs, 2025), we further generate an SFT dataset for post-training by leveraging the questions in these triplets, using DeepSeek R1 (DeepSeek-AI et al., 2025) as the response generator to generate Chain-of-Thought responses. To ensure the quality of the generated responses, we generate 3 times for each question and perform test-based reject sampling, yielding a large-scale, high-quality, and verifiably correct SFT dataset for coding. We refer to this dataset as KodCode-SFT.
@@ 3.1 Pipeline Analysis
Effectiveness of Self-Verification. To evaluate the reliability of our self-verification pipeline, we conduct experiments using the MBPP validation dataset (Austin et al., 2021), which contains 90 coding questions with ground-truth unit tests that were manually written and verified by humans. Following our pipeline, we generate both solutions and unit tests for these 90 questions. 10 questions are discarded as they fail all self-verification attempts, leaving us with 80 solutions that pass self-verification. We then evaluate these retained solutions against the MBPP ground-truth unit tests. We note that due to ambiguity in MBPP questions, some solutions that follow correct logic fail assertions due to mismatched input formats or numerical precision differences. After manual review of these edge cases, 78 out of 80 solutions pass all ground-truth unit tests (97.5% pass rate). This high success rate demonstrates our pipeline’s effectiveness in generating verified solutions. Please refer to Appendix B.1 for analysis of the two failure cases.
@@
Figure 2: Statistics on pass rates via self-verification in Step 2 by subset with varing number of attempts.
@@ 3.2 Dataset Statistics and Analysis
Figure 3: Distribution of token lengths for questions and responses, along with unit test counts across different subsets.
@@
The visualization in Figure 4 reveals two key observations. First, KodCode’s question distribution (shown in color) spans the entire space, while baseline datasets (in gray) cluster primarily in the upper left region, demonstrating KodCode’s broader topical diversity. Second, the Algorithm and Filter subsets of KodCode show comprehensive coverage across the entire space, validating their role in enhancing KodCode’s overall diversity and robustness.
Difficulty. We analyze the difficulty distribution across KodCode  subsets by examining the success rate of $(sol,test)$ pairs in self-verification across $n=10$ attempts per question, as shown in Figure 5. We categorize questions into four difficulty levels: easy (pass rate >2/3), medium (1/3 to 2/3), hard (<1/3), and fail (all failures). The analysis reveals that Prefill tasks are typically the easiest, while Codeforces, Taco, and Code Contests subsets present the greatest challenges, as shown in their higher failure rates and larger proportions of hard questions.
@@
KodCode KodCode-32B-50K 92.7 85.4 89.9 76.2 59.8 37.8 51.1 32.4 87.8 35.9 6.7 61.22
KodCode-32B-18K-Hard 90.9 86.6 89.2 77.0 59.7 37.2 50.5 31.1 90.7 39.3 5.6 61.26
@@ 4 Performance Evaluation
To evaluate the performance of the KodCode  dataset, we conduct supervised fine-tuning using Qwen2.5-Coder-32B-Instruct Hui et al. (2024) on the KodCode-SFT dataset. We then compare the model’s performance across several widely used code generation benchmarks against existing model baselines.
@@ 4.1 Experimental Setup
KodCode  Setup. We evaluate two variants of KodCode-SFT. For preprocessing, we exclude R1 responses that are too long, too short, or implement class-based solutions rather than functional implementations. The first version, KodCode-50K, contains 50K instruction-response pairs selected based on the empirical results of mixing data from different sources and difficulties. The second version, KodCode-Hard-18K, contains all 18K instructions that are labeled as hard after applying the same preprocessing filters. We refer to the models fine-tuned on these two datasets as KodCode-32B-50K, and KodCode-32B-18K-Hard.
For SFT, we use a cosine learning rate schedule with a maximum learning rate of $1\times 10^{-5}$ when fine-tuning the Qwen-2.5 model. The maximum sequence length is 16384. The detailed training configurations can be found in Appendix C.1.
Baselines. We compare our KodCode-tuned model against several strong baselines, including non-reasoning models (Llama-3.3-70B-Instruct (Dubey et al., 2024), Llama-3.1-Tulu-3-70B (Lambert et al., 2025), Qwen-2.5-32B/72B-Instruct (Team, 2024a), Qwen-2.5-Coder-32B-Instruct (Hui et al., 2024)), and reasoning models (QwQ-32B-Preview (Team, 2024b), DeepSeek-R1-Distill-Llama-70B (DeepSeek-AI et al., 2025), Sky-T1-32B-Preview (Team, 2025), and Bespoke-Stratos-32B (Labs, 2025)).
@@ 4.2 Experimental Results
Table 2 presents a comparative analysis of Qwen-2.5-32B-Coder-Instruct fine-tuned on KodCode-SFT against various baseline models. Our fine-tuned models consistently outperform all baselines, including larger models, across both the Complete and Instruct subsets of BigCodeBench. Specifically, for BigCodeBench-C, our model achieves 59.8% (+1.9%) in the Full category and 37.8% (+4.7%) in the Hard category, compared to the strongest baseline. For BigCodeBench-I, it achieves 51.1% (+1.7%) in the Full category and 32.4% (+5.4%) in the Hard category, demonstrating strong performance across different evaluation settings. Additionally, on HumanEval, our model reaches 92.7%, surpassing Qwen2.5-Coder-32B-Instruct (90.9%) by 1.8%. For LiveCodeBench-Easy, our model attains 90.7%, exceeding all baselines, including QwQ-32B-Preview (90.0%). Overall, our model achieves an average score of 61.26%, the highest among all evaluated models, highlighting the quality and diversity of our KodCode dataset and its effectiveness in improving code generation performance across various benchmarks.
@@
To validate the impact of challenging instances in KodCode, we compare models trained on two different 10K sample sets: one randomly sampled from the KodCode-50K dataset (named as KodCode-10K) and another from the KodCode-Hard-18K dataset (named as KodCode-Hard-10K). As shown in Table 3, KodCode-Hard-10K outperforms KodCode-10K on BigCodeBench-I Hard (31.8% vs. 27.7%, +4.1%) and BigCodeBench-I Full (50.6% vs. 49.9%, +0.7%), confirming that exposure to difficult coding problems improves model robustness. Similarly, for BigCodeBench-C Hard, KodCode-Hard-10K achieves 39.9%, surpassing KodCode-10K (38.5%, +1.4%). On LiveCodeBench Hard, KodCode-Hard 10K leads (6.3% vs. 4.8%, +1.5%), reinforcing that hard samples do enhance performance on complex programming tasks.
Benchmarks KodCode Hard-10K KodCode -10K KodCode NoConvert-10K
BigCodeBench-C Full (%) 60.4 61.1 60.3
BigCodeBench-C Hard (%) 39.9 38.5 35.1
BigCodeBench-I Full (%) 50.6 49.9 49.6
BigCodeBench-I Hard (%) 31.8 27.7 28.4
@@
To assess the impact of the style converter, we remove all instances processed by the style converter from the KodCode-50K dataset and randomly sample 10K instances from the remaining datasets for fine-tuning, naming the resulting dataset KodCode-NoConvert-10K. As shown in Table 3, KodCode-NoConvert-10K performs lower on BigCodeBench-C Full (60.3% vs. 61.1%) and BigCodeBench-C Hard (35.1% vs. 38.5%), indicating that removing style variations slightly reduces performance. In LiveCodeBench Easy, KodCode-NoConvert-10K scores 86.4%, lower than KodCode-10K (87.8%). This result is consistent with findings from (Li et al., 2025), highlighting that question format plays a role in code LLM performance.
@@ 6 Conclusion and Future Work
While models fine-tuned on KodCode achieve state-of-the-art performance across most benchmarks, their performance on LiveCodeBench-Hard remains limited. This gap likely stems from insufficient representation of highly challenging competition-level programming problems in our dataset, which are prevalent in LiveCodeBench-Hard. Our current approach is also limited to using SFT to validate the quality of question–solution–test triplets. Future work could explore reinforcement learning techniques such as PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), and REINFORCE++ (Hu, 2025) to further enhance model performance.
@@ C.1 Supervised Fine-Tuning Setups
We adapt the system prompt from Team (2025), modifying it to better accommodate R1-style long chain-of-thought responses during fine-tuning. The full prompt can be found in Figure 13.

### v2

@@ 1 Introduction
Ideally, training high-performing coding LLMs requires high-quality data with verified solutions and test cases for post-training stages, including supervised fine-tuning (SFT) and reinforcement learning (RL) DeepSeek-AI et al. (2025); Team (2025a); Hui et al. (2024); Wei et al. (2024a). While human-curated coding datasets like TACO Li et al. (2023), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022) offer high quality questions, canonical solutions, and tests, their limited scale constrains model training. Synthetic datasets have emerged as an alternative Wang et al. (2023); Long et al. (2024), but often lack diversity Xu et al. (2024), sufficient complexity Luo et al. (2023), and reliable response verification Lei et al. (2024).
@@
Furthermore, to validate KodCode’s effectiveness for code LLM post-training, we evaluate models using supervised SFT and RL across standard benchmarks: HumanEval(+), MBPP(+), BigCodeBench, and LiveCodeBench. Our experimental results show that KodCode-fine-tuned models achieve state-of-the-art performance, surpassing other open-source models such as Qwen2.5-Coder-32B-Instruct and DeepSeek-R1-Distill-Llama-70B in most of the benchmarks.
@@ 2.2 Step 2: Solution & Test Generation
To generate verifiably correct coding solutions and unit tests $(sol,test)$ for questions from Step 1, we employ a self-verification procedure as detailed below. To ensure quality of solution and tests, we first employ GPT-4o-0513 (which achieves state-of-the-art performance among non-reasoning models on the BigCodeBench Leaderboard (Zhuo et al., 2024)) to generate both solution and test, then execute these unit tests to validate the correctness of the solution. Additionally, we perform branch coverage analysis using the pytest-cov framework to ensure the diversity of test cases. Only question-solution-test triplets that pass self-verification and achieve 100% branch coverage are retained. The prompt template is provided in Appendix D.6.
Since our goal is to generate verifiable training data that is challenging and diverse in coverage, a key challenge arises: even state-of-the-art models cannot guarantee bug-free code or ensure that solutions pass their unit tests. Simply discarding questions that fail self-verification risks eliminating many challenging ones, potentially biasing our dataset towards a distribution of simple problems.
Our solution to address this challenge is to assign additional self-verification attempts for hard questions. For each question from Step 1, we allow up to a maximum of $n$ attempts (where $n=10$ in our experiments) to generate a solution that passes its unit tests. Importantly, each attempt regenerates both the solution and its corresponding unit tests from scratch, since if the initial tests contain errors, all subsequent solutions would fail regardless of their correctness. To maintain the quality of regenerated test cases, we only retain new $(sol,test)$ that achieve equal or higher branch coverage compared to previous attempts. This ensures that regenerated tests do not become progressively simpler or less comprehensive.
We note that our approach can preserve challenging questions while naturally assigning difficulty labels based on the success rate across attempts. Questions that fail to generate correct solutions after $n$ attempts are discarded as they likely contain inherent flaws. Upon completing Step 2, we obtain a collection of 279K verified triplets.
@@ 2.3 Step 3: Post-training Data Synthesis
Motivated by recent advances in reasoning models (Team, 2025a; Labs, 2025), we further generate an SFT dataset for post-training by leveraging the questions in these triplets, using DeepSeek R1 (DeepSeek-AI et al., 2025) as the response generator to generate Chain-of-Thought responses. To ensure the quality of the generated responses, we generate 3 times for each question and perform test-based reject sampling, yielding a large-scale, high-quality, and verifiably correct SFT dataset for coding. We refer to this dataset as KodCode-SFT.
@@ 3.1 Pipeline Analysis
Effectiveness of Self-Verification. To evaluate the reliability of our self-verification pipeline, we conduct experiments on the MBPP validation dataset (Austin et al., 2021) and the LiveCodeBench-V5 test set (Jain et al., 2024). For MBPP, we utilize all 90 coding questions that include ground-truth unit tests manually written and verified by humans. For LiveCodeBench-V5, we select all 381 questions that provide starter code and functional test cases.
Following our pipeline, we generate both solutions and unit tests for these questions. After applying our Self-Verification approach, 80 solutions (88.9%) pass self-verification for MBPP, while 190 solutions (49.9%) pass self-verification for LiveCodeBench-V5. We then evaluate these retained solutions against the ground-truth unit tests.
We note that due to inherent ambiguity in MBPP questions, some solutions that follow correct logic fail assertions because of mismatched input formats or numerical precision differences. After manual review of these edge cases, 78 out of 80 solutions pass all ground-truth unit tests, achieving a 97.5% pass rate for MBPP. For LiveCodeBench-V5, 189 solutions (99.47%) successfully pass all tests. This high success rate demonstrates the effectiveness of our pipeline in generating verified solutions. Please refer to Appendix B.1 for detailed analysis of the failure cases.
@@
Figure 2: Statistics on pass rates via self-verification in Step 2 by subset with varying number of attempts.
Figure 3: Distribution of token lengths for questions and responses, along with unit test counts across subsets.
@@ 3.2 Dataset Statistics and Analysis
The visualization in Figure 4 reveals two key observations. First, KodCode’s question distribution (shown in color) spans the entire space, while baseline datasets (in gray) cluster primarily in the upper left region, demonstrating KodCode’s broader topical diversity. Second, the Algorithm and Filter subsets of KodCode show comprehensive coverage across the entire space, validating their role in enhancing KodCode’s overall diversity.
Difficulty. We analyze the difficulty distribution across KodCode  subsets by examining the success rate of $(sol,test)$ pairs in self-verification across $n=10$ attempts per question, as shown in Figure 5. We categorize questions into four difficulty levels: easy (pass rate >2/3), medium (1/3 to 2/3), hard (<1/3), and fail (all failures). The analysis reveals that the Prefill subset are typically the easiest, while Codeforces, Taco, and Code Contests subsets are more challenging, as shown in their higher failure rates and larger proportions of hard questions.
@@
KodCode KodCode-32B-SFT-50K 92.7 85.4 89.9 76.2 59.8 37.8 51.1 32.4 87.8 35.9 6.7 61.22
KodCode-32B-SFT-Hard-18K 90.9 86.6 89.2 77.0 59.7 37.2 50.5 31.1 90.7 39.3 5.6 61.26
@@ 4 Performance Evaluation
To evaluate the performance of the KodCode  dataset, we conduct both supervised fine-tuning using Qwen2.5-Coder-32B-Instruct Hui et al. (2024) on the KodCode-SFT dataset, and GRPO Shao et al. (2024) using Qwen2.5-7B-Instruct-1M Team (2025b) and Qwen2.5-Coder-7B-Instruct Hui et al. (2024) on the KodCode dataset. We compare the model’s performance across several widely used code generation benchmarks against existing model baselines.
@@ 4.1 Experimental Setup
SFT Setup. We evaluate two variants of KodCode-SFT. For preprocessing, we exclude R1 responses that are too long, too short, or implement class-based solutions rather than functional implementations. The first version, KodCode-SFT-50K, contains 50K instruction-response pairs selected based on the empirical results of mixing data from different sources and difficulties. The second version, KodCode-SFT-Hard-18K, contains all 18K instructions that are labeled as hard after applying the same preprocessing filters. We refer to the models fine-tuned on these two datasets as KodCode-32B-SFT-50K, and KodCode-32B-SFT-Hard-18K. We use a cosine learning rate schedule with a maximum learning rate of $1\times 10^{-5}$ when fine-tuning the Qwen-2.5 model. The maximum sequence length is 16384. The detailed training configurations can be found in Appendix C.1.
RL Setup. We conduct RL experiments on both Qwen2.5-7B-Instruct-1M Team (2025b) and Qwen2.5-Coder-7B-Instruct Hui et al. (2024) using 9.5K randomly selected samples from KodCodeas the training set, and 0.5K as the validation set. We refer to this dataset as KodCode -RL-10K. We perform GRPO Shao et al. (2024) for 256 steps with actor learning rate of $5\times 10^{-7}$, 16 rollouts per question, a batch size of 256, max response length of 4096, and apply KL coefficient of 0.001. We assign binary rewards for RL: 1 if the model’s solution passes all unit tests, and 0 otherwise.
Baselines. We compare our KodCode-tuned model against several strong baselines, including non-reasoning models (Llama-3.3-70B-Instruct (Dubey et al., 2024), Llama-3.1-Tulu-3-70B (Lambert et al., 2025), Qwen-2.5-32B/72B-Instruct (Team, 2024a), Qwen-2.5-Coder-32B-Instruct (Hui et al., 2024)), and reasoning models (QwQ-32B-Preview (Team, 2024b), DeepSeek-R1-Distill-Llama-70B (DeepSeek-AI et al., 2025), Sky-T1-32B-Preview (Team, 2025a), and Bespoke-Stratos-32B (Labs, 2025)).
@@ 4.2 Experimental Results
Table 2 presents a comparative analysis of Qwen-2.5-32B-Coder-Instruct fine-tuned on KodCode-SFT against various baseline models. Our fine-tuned SFT models consistently outperform all baselines, including larger models, across both the Complete and Instruct subsets of BigCodeBench. Specifically, for BigCodeBench-C, our model achieves 59.8% (+1.9%) in the Full category and 37.8% (+4.7%) in the Hard category, compared to the strongest baseline. For BigCodeBench-I, it achieves 51.1% (+1.7%) in the Full category and 32.4% (+5.4%) in the Hard category, demonstrating strong performance across different evaluation settings. Additionally, on HumanEval, our model reaches 92.7%, surpassing Qwen2.5-Coder-32B-Instruct (90.9%) by 1.8%. For LiveCodeBench-Easy, our model attains 90.7%, exceeding all baselines, including QwQ-32B-Preview (90.0%). Overall, our model achieves an average score of 61.26%, the highest among all evaluated models, highlighting the quality and diversity of our KodCode dataset and its effectiveness in improving code generation performance across various benchmarks.
@@
To validate the impact of challenging instances in KodCode, we compare models trained on two different 10K sample sets: one randomly sampled from the KodCode-SFT-50K dataset (named as KodCode-SFT-10K) and another from the KodCode-SFT-Hard-18K dataset (named as KodCode-SFT-Hard-10K). As shown in Table 3, KodCode-SFT-Hard-10K outperforms KodCode-SFT-10K on BigCodeBench-I Hard (31.8% vs. 27.7%, +4.1%) and BigCodeBench-I Full (50.6% vs. 49.9%, +0.7%), confirming that exposure to difficult coding problems improves model robustness. Similarly, for BigCodeBench-C Hard, KodCode-SFT-Hard-10K achieves 39.9%, surpassing KodCode-SFT-10K (38.5%, +1.4%). On LiveCodeBench Hard, KodCode-SFT-Hard-10K leads (6.3% vs. 4.8%, +1.5%), reinforcing that hard samples do enhance performance on complex programming tasks.
Benchmarks KodCode SFT-Hard-10K KodCode SFT-10K KodCode-SFT NoConvert-10K
BigCodeBench-C Full 60.4 61.1 60.3
BigCodeBench-C Hard 39.9 38.5 35.1
BigCodeBench-I Full 50.6 49.9 49.6
BigCodeBench-I Hard 31.8 27.7 28.4
@@
To assess the impact of the style converter, we remove all instances processed by the style converter from the KodCode-SFT-50K dataset and randomly sample 10K instances from the remaining datasets for fine-tuning, naming the resulting dataset KodCode-SFT-NoConvert-10K. As shown in Table 3, KodCode-SFT-NoConvert-10K performs lower on BigCodeBench-C Full (60.3% vs. 61.1%) and BigCodeBench-C Hard (35.1% vs. 38.5%), indicating that removing style variations slightly reduces performance. In LiveCodeBench Easy, KodCode-SFT-NoConvert-10K scores 86.4%, lower than KodCode-SFT-10K (87.8%). This result is consistent with the findings of (Li et al., 2025), highlighting that the question format plays a role in the performance of the LLM code.
Effectiveness of KodCode on RL.
Model LiveCodeBench BCB-Complete BCB-Instruct HumanEval MBPP Average
Easy Medium Hard Full Hard Full Hard Base Plus Base Plus
Qwen2.5-Coder-7B-Instruct 57.4 23.0 4.4 52.0 21.6 41.8 19.6 91.5 85.4 83.1 71.7 52.32
+ RL KodCode-10K (Step 128) 65.2 21.1 4.1 52.5 25.7 42.2 20.3 90.9 86.0 84.9 72.8 53.56
+ RL KodCode-10K (Step 256) 64.5 19.9 3.3 53.7 27.0 42.9 21.6 90.2 85.4 86.5 74.1 53.99
Qwen2.5-7B-Instruct-1M 57.7 12.4 3.7 45.3 14.2 36.6 17.6 86.0 79.3 78.8 69.3 47.63
+ RL KodCode-10K (Step 128) 60.2 19.0 2.6 47.0 19.6 36.7 13.5 90.2 83.5 81.0 70.9 49.69
+ RL KodCode-10K (Step 256) 57.0 18.7 3.0 48.2 19.6 36.8 12.8 91.5 86.0 82.8 72.8 50.30
Table 4: This table evaluates the performance of models before and after training with GRPO using KodCode-RL-10K. Notable performance gains are observed across most benchmarks compared to the baselines.
Table 4 compares the performance of models before and after GRPO using KodCode-RL-10K. We observe significant performance improvement on most of the benchmarks compared to the original model after RL. In addition, we observe that continuing to increase the training steps can further enhance the model performance. This indicates the effectiveness of KodCode for RL training.
@@ 6 Conclusion and Future Work
Acknowledgment
This work is partially supported by the Office of Naval Research (ONR) under grant N0014-23-1-2386, the Air Force Office of Scientific Research (AFOSR) under grant FA9550-23-1-0208, and the National Science Foundation (NSF) AI Institute for Agent-based Cyber Threat Intelligence and Operation (ACTION) under grant IIS 2229876.
This work is supported in part by funds provided by the National Science Foundation, Department of Homeland Security, and IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF or its federal agency and industry partners.
@@ Acknowledgment
While models fine-tuned on KodCode achieve state-of-the-art performance across most coding benchmarks, their performance on LiveCodeBench-Hard remains limited. This gap likely stems from insufficient representation of highly challenging competition-level programming problems in our dataset, which are prevalent in LiveCodeBench-Hard. Future work could explore methods to synthesis highly challenging coding problems.
@@ C.1 Supervised Fine-Tuning Setups
We adapt the system prompt from Team (2025a), modifying it to better accommodate R1-style long chain-of-thought responses during fine-tuning. The full prompt can be found in Figure 13.